Tohban Report 2013-01-02

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 26 Dec 2012
End Date: 2 Jan 2013
Tohban: Pascal Saint-Hilaire
Tohban email:
Next Tohban: Juan-Carlos Martinez Oliveros
List all reports


Solar Activity

A few C-flares over Christmas, then pretty quiet. Two C-flares in the last few days. These were all produced by AR 11635, which has now rotated off the visible solar disc. There are a few current and a couple of new active regions, but they have not produced anything above a B-class flare so far. STEREO-B images show that a new ARs should appear later in the (Tohban) week.

How many GOES flares occurred?

 Flares above B, C, M, X class were     22    13     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     12     9     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      9     7     0     0

There were RHESSI flares/GOES flares 185 / 35 over the time range 25-Dec-12 01-Jan-13

Memory Management

The SSR fill level has been regularly reaching 0% 1 or 2 passes before the end of the daily pass set. Hence, we plan on switching to the Normal/Vigorous decimation scheme this week, and after a day or two, to switch on rear segments at night.

Spacecraft Status

Cold Plate 2 at 107 K. More or less as expected. We are considering playing with D2 voltages in a few weeks, once the daily pass set occurs during the day, in the hope it'll segment.

Data Gaps

No data gaps of note this week. However, the large data gap reported last week: " There was a large data gap due to the Wallop station problem.

 GAP START TIME              GAP END TIME                   GAP (SEC)
 2012-12-20T01:10:00.000 -- 2012-12-20T04:55:00.000       13500.000

" still has no Observing Summary data generated for it. The Level-0 orbit files corresponding to these times indeed do not have proper sizes.

Jim says the VC files are corrupted and is looking into it.

Detector issues

Basically, no issues!

Detector 8: From Jeremy: "I turned on detector 8 events were finally turned on again around 2012-363-19:00 UTC. On the next orbit RHESSI flew through the SAA and there was no significant increase in the SSR fill level. We put in about 0.1 % more of the total SSR than the previous orbit, although the SAA passage ended in eclipse."

Since then, detector 8 has been very well behaved, even after SAA passage. The number of front events is perhaps 20-30% above that of other detectors (1,3,4,6,7,9), but not as "bad" as detector 5 (about 50% more events).

Detector 4: Could not find a single de-segmentation event from Dec 26 to Jan 1 or so! I believe it has been nursed back to full health.

Some other issues:

The GOES(or EVE proxy)+RHESSI plots usually accessible from the Browser appear to not being made anymore, since 2012/12/24. Will notify Albert.

Decimation active/vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 min
Night time data (rears) +/- 4 min
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? None
Personal tools